A 50 kg woman sits near the edge of a 10 kg table (table: 4m in length, legs at 1m & 3m, she sits at 0m). If the table does not tilt and fall over what is the MINIMUM distance from the table's edge where she can sit?
Answer #1

Applying equilibrium of moment

Here, Ftable * (L/2 - 1) = Fwoman * (1 - x)

=> 10 * 9.8 * (4/2 - 1) = 50 * 9.8 * (1 - x)

=> x =   0.8 m

Thus, the minimum distance from the table's edge where she can sit = 0.8 m